# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ III.4.2 Single and two-particle production in $e+A$ and $p+A$
Figure 10: The nuclear modification factor of forward charged hadron production in $p+{\rm Pb}$ collisions as a function of $p_{T}$ (top) and rapidity (bottom) at $\sqrt{s_{NN}}=5$ TeV from the EPPS21 nPDFs at a factorization scale of $\mu={\rm max}\{m_{c},p_{T}\}$ compared to NLO CGC calculations including high energy, Sudakov and threshold resummation [321]. The blue band shows the nPDF uncertainty while the red band shows the CGC uncertainties arising from varying the factorization scale, assuming the $p+\mathrm{Pb}$ and $p+p$ cross sections are fully correlated. The calculations are compared to LHCb charged hadron data [322]. The corresponding $\pi^{0}$ data at $\sqrt{s_{NN}}=8.16\,{\rm TeV}$ [323] are shown for comparison as a function of $p_{T}$. Note the large nPDF uncertainty at low $p_{T}$.
@@
The saturation and nPDF pictures also lead to distinct predictions of single-inclusive hadron production. Here, the features of inclusive hadron production in $p{\rm Pb}$ collisions employing the CGC and nPDF approaches are compared.
Figure 10 compares the nuclear modification factors in the CGC [321] and collinear factorization approaches as functions of $p_{\rm T}$ and rapidity. The CGC uncertainty bands include variations of the factorization scale and the “auxiliary” scale assuming the $p+\mathrm{Pb}$ and $p+p$ cross sections are correlated. The uncertainty bands in the nPDF calculations reflect the variation of the nPDF parameters added in quadrature. This variation is a reflection of the precision of the data used to extract the nPDFs. The precision of the data is the leading uncertainty above $p_{T}\gtrsim 3$ GeV. At lower $p_{T}$ the scale uncertainties becomes presumably very large but are not well defined because the factorization scale becomes smaller than the nPDF parametrization scale.
In general, the CGC predictions for single-inclusive hadrons tend to have somewhat steeper slopes in $p_{T}$ and $y$ compared to the nPDF predictions. The trend for charged hadrons in this picture is in somewhat better agreement with the LHCb data [322]. However, the baryon contribution to charged hadron production results in a poorer description of charged hadron production below $p_{T}\sim 10$ GeV using standard independent fragmentation functions [316]. The discrepancy with the is probably most noticable in $p/\pi$ data which exhibits a bump at $p_{T}\sim 3$ GeV [324] which cannot be described by modifying the fragmentation functions. Indeed, the pion $p_{\rm T}$ slope is flatter [323] and better in line with the nPDF calculations.
Because the CGC and nPDF calculations both depend on the fragmentation functions, forward pion measurements differential in $p_{T}$ and $y$ would be more beneficial than inclusive charged hadrons alone. Based on Fig. 10, the nPDF and CGC calculations could be distinguished by systematic studies of the $p_{T}$ and rapidity slopes at $p_{T}\gtrsim 3$ GeV. In $D$ meson studies this $p_{T}$ restriction could be relaxed because the finite charm quark mass allows the perturbative calculation to be continued to $p_{T}=0$. Existing LHCb measurements in the forward region [124, 127] could be studied from this perspective.
@@
Single inclusive hadron production (SIDIS), where a hadron with transverse momentum $p_{T}$ and longitudinal momentum fraction $z=(p_{h}\cdot P_{N})/(q\cdot P_{N})$ is measured in the final state, could also distinguish between saturation and shadowing effects.
@@
Future SIDIS measurements at the EIC are known [325] to significantly reduce the uncertainties in the PDFs in $e+p$ collisions, particularly for radiative production of light sea quarks. Analogous measurements in $e+A$ collisions should therefore provide valuable information about sea quark relative to gluon shadowing in the nPDFs.
@@
At small $x$, the dominant parton-level channel for SIDIS is $\gamma^{*}g\to q\bar{q}$ where the kinematics of one of the final state quarks is integrated over. SIDIS has the advantage of probing smaller values of $x$ for $t$-channel gluons compared to dihadron production, discussed in the previous section, because the gluon $x$ is comparable to Bjorken $x$ for $p_{T}\ll Q$.
@@
In fact, when $p_{T}\ll Q$, SIDIS also exhibits TMD factorization [326, 327, 292, 328] for the sea quarks: the CGC cross section can be expressed [329] as product of a hard factor for the $\gamma^{*}q\to q$ subprocess at scale $Q$ and a TMD sea quark distribution [330, 331, 332] resulting from $g\to q\overline{q}$ splitting in the nuclear target with the $q$ or $\overline{q}$ measured at final-state transverse momentum $p_{T}$. SIDIS can thus probe saturated sea quarks in the target wave function with $p_{T}\sim Q_{s}\ll Q$ [329]. An additional interesting regime of gluon saturation is the very forward limit [333], where $1-z\ll 1$, such that $(1-z)Q^{2}\sim Q_{s}^{2}$ with $Q\gg Q_{s}$. In this regime, the CGC predicts a strong suppression of $R_{eA}$ as a function of $z$ as $z\to 1$ [333].
@@
Due to the hierarchy of scales imposed in these situations, radiative corrections are dominated by Sudakov logarithms, as in two-particle correlations, as previously discussed. The all-order exponentiation of these logarithms will also suppress the cross section, competing with the suppression caused by saturation. As in the case of two-particle correlations, these corrections make the physical interpretation of the suppression more difficult. Theoretical frameworks, including both suppression mechanisms, either within the TMD factorization formalism [295] or by coupling the CGC to a final-state parton shower [334, 301], are the only way to conclusively determine the role of nonlinear effects in future experimental data.
@@
In the absence of a specific hierarchy of scales imposed on the final state, the Sudakov effects disappear. However, the interesting regime for saturation is typically the semi-hard one where $p_{T}\sim Q\sim Q_{s}$ and $z\sim 1/2$. In this regime, no approximation is applicable and the all-twist CGC cross section must be employed. While LO estimates are simple to obtain because the SIDIS cross-section involves only dipoles, NLO predictions are challenging to compute numerically due to the complexity of the formulas [199, 195, 200]. A direct comparison to the leading-twist nPDF approach necessitates at least NLO precision and more developments are needed to evaluate CGC cross sections at full NLO.
@@
Single-inclusive jet production [335, 336, 337, 338], which has less nonperturbative model dependence in the final state because no fragmentation functions are required, is related to SIDIS. At low $p_{T}$, $p_{T}\ll Q$, jets can probe the sea quark distribution, much like single inclusive hadron production [339]. Recently, jet distance measures have been developed to ensure TMD factorization beyond leading order using jets [340, 339]. However, jet measurements at low $p_{T}$ remain challenging [341]. (At HERA, jets have been measured down to $4-5$ GeV in the Breit frame.) In addition, the phase space for low $p_{T}$ jets at high $Q^{2}$ and small-$x$ shrinks drastically in EIC kinematics. At moderate $x$, single-inclusive jet production at low $p_{T}$ is also sensitive to jet energy loss in cold nuclear matter [342], see the discussion in the following section. Lastly, it is also possible to study saturation effects in jet SIDIS in the target fragmentation region of DIS in the Breit frame [343]. Furthermore, nucleon energy correlators have been proposed as a new way to study saturation dynamics in this kinematic regime [344].
@@ III.4.3 Diffractive vector meson production in $\gamma+A$
Figure 11: A compilation of exclusive $J/\psi$ production measurements in Pb+Pb UPCs at $\sqrt{s_{NN}}=5.02\,{\rm GeV}$ [135, 136, 345, 139, 346]. (Top) The curves represent NLO calculations with EPPS21 nPDFs using different choices for the factorization and renormalization scales $\mu$. The light blue band is the EPPS21 uncertainty at $\mu=2.39$ GeV. The NLO calculations are from Refs. [347, 348]. (Bottom) The calculations here, taken from Ref. [349], representing different realizations of saturation-based calculations, depending on their description of the dipole scattering with the nucleus, are based on [350, 351, 352, 353, 354, 355].
@@
Diffractive vector meson production in photon-nucleus, $\gamma+A$, interactions serves as a complementary and particularly sensitive tool to probe gluon structure at small $x$ [356], relative to other observables discussed earlier. Unlike semi-inclusive observables, diffraction demands a color neutral exchange, which at LO is achieved by at least two-gluon exchange, making it more sensitive to saturation and shadowing effects. There are two kinds of reactions: coherent, where the nucleus remains intact and interactions access an average gluon distribution; and incoherent reactions, characterized by nuclear breakup and sensitive to event-by-event fluctuations of the gluon densities [357, 358, 359, 356]. Diffractive $J/\psi$ production is particularly compelling because the charm mass provides a hard scale, allowing for perturbative treatment, but also comparable to the saturation scale in $\gamma+A$ collisions.
@@
In the saturation formalism, this process is studied within the dipole picture, where, at LO, the photon splits into a quark-antiquark $q\overline{q}$ pair [360] (and $q\overline{q}g$ at next-to-leading order [187]) which interacts with the nucleus through multiple coherent scatterings and then recombines into the vector meson. The saturation physics is encoded in the $q\overline{q}$ with the nucleus, characterized by the dipole amplitude. The energy dependence of the dipole is governed by the BFKL equation in the linear regime and the BK/JIMWLK in the nonlinear regime.
@@
The cross section for diffractive vector meson production in $\gamma+A$ interactions has been studied as a function of the center of mass energy and rapidity of the produced vector meson in ultraperipheral collisions of heavy-ions at RHIC and the LHC. In the linear regime, governed by BFKL evolution, the cross section is predicted to grow as a power of the energy while saturation effects are expected to suppress this growth. The suppression is expected to be strong when the typical transverse size of the vector meson is larger than or comparable to the inverse of the saturation scale [361]. At the energies and rapidities reached in these colliders, the growth of the $J/\psi$ cross section in $\gamma+p$ collisions in UPCs is consistent with a power law because the typical transverse size of $J/\psi$ is smaller than the inverse of the saturation scale in the proton [362, 363]. On the other hand, strong nuclear suppression is seen in $J/\psi$ production in $\gamma+{\rm Pb}$ collisions at the LHC. The suppression is stronger than predicted by the current saturation-based approaches, see Fig. 11. These different saturation-based predictions differ in their description of the dipole amplitude but are usually constrained by fits to $\gamma+p$ data. The CGC (MSS) calculation starts with initial conditions taken from an impact-parameter dependent MV model with the rapidity and energy evolution following the JIMWLK equation [350], while the b-BK model (BCCM) employs an impact-parameter dependence dipole amplitude that evolves according to impact-parameter dependent BK evolution [351]. The BGK-I (LM) [352], IPsat (LM) [353], GG-HS (CCK) [354], and IIM BG (GM) [355] calculations implement different phenomenological models of the impact parameter, rapidity and dipole size dependence of the dipole amplitude. For more details on the implementation of the different models see Ref. [349].
@@
Exclusive $J/\psi$ production in Pb+Pb collisions probes the nPDFs predominantly at $10^{-5}\lesssim x\lesssim 10^{-2}$ and scales similar to those in low $p_{T}$ open charm production [347, 348]. The coefficient functions are known up to NLO in strong coupling [364] when the $J/\psi$ is treated in nonrelativistically. The extraction of nuclear PDFs, or testing their dynamical modeling as in e.g. leading-twist shadowing [148, 365], based on UPC data is complicated by several factors. Convergence of the perturbative expansion has proven to be quite unstable due to coincidental cancellations between the LO and NLO gluon contributions. As a result, the calculated cross sections are rather sensitive to the factorization and renormalization scales, see Fig. 11. Better stability seems to be achieved by including resummation of BFKL-type terms [366]. Other complications arise because generalized nuclear parton distributions enter the calculations. These GPDs can be approximated by nuclear PDFs but the associated theoretical uncertainty is difficult to reliably assess. It at least induces some degree of model dependence. Other theoretical uncertainties include the role of nonrelativistic corrections to the treatment of the $J/\psi$ wave function as well as theoretical modeling of the exclusivity condition. All in all, an accurate description of exclusive $J/\psi$ at Pb+Pb UPCs is still a distant goal which hinders the distinction of dynamical effects of saturation. Currently there also appears to be tensions between the ALICE [135] and preliminary ATLAS [346] data, see Fig. 11.
@@
Looking forward, studying the energy dependence across different colliding systems, such as $\gamma+{\rm O}$, as well as the production of lighter and heavier vector mesons such as $\phi$ and $\Upsilon$ could help further discriminate different approaches to the observed nuclear suppression and distinguish between saturation and nonperturbative shadowing. Moreover, the production ratios of excited vector meson states to the ground states, particularly $\psi$(2S) to $J/\psi$, offers additional discriminatory power. In Ref. [367] it was argued that the distinctive node in the $\psi$(2S) wavefunction enhances the sensitivity to nonlinear effects, providing a promising ameans to discriminate between linear shadowing and nonlinear saturation dynamics. Lastly, further measurements of incoherent reactions, where the nucleus breaks up, and which are sensitive to fluctuations, could also shed some light on small-$x$ dynamics [368].

### v2

@@ III.4.2 Single and two-particle production in $e+A$ and $p+A$
Figure 10: The nuclear modification factor of forward charged hadron production in $p+{\rm Pb}$ collisions as a function of $p_{T}$ (top) and $y$ (bottom) at $\sqrt{s_{NN}}=5$ TeV from the EPPS21 nPDFs at factorization scale $\mu={\rm max}\{m_{c},p_{T}\}$ compared to NLO CGC calculations including high energy, Sudakov and threshold resummation [323]. The blue band shows the nPDF uncertainty while the red band shows the CGC uncertainties arising from varying the factorization scale, assuming the $p+\mathrm{Pb}$ and $p+p$ cross sections are fully correlated. The calculations are compared to LHCb charged hadron data [324]. The corresponding $\pi^{0}$ data at $\sqrt{s_{NN}}=8.16\,{\rm TeV}$ [325] are shown for comparison as a function of $p_{T}$. Note the large nPDF uncertainty at low $p_{T}$.
@@
The saturation and nPDF approaches also lead to distinct predictions of single-inclusive hadron production. Here, the features of inclusive hadron production in $p+{\rm Pb}$ collisions employing the CGC and nPDF approaches are compared.
Figure 10 compares the nuclear modification factors in the CGC [323] and collinear factorization approaches as functions of $p_{\rm T}$ and $y$. The CGC uncertainty bands are dominated by variations of the factorization scale, assuming the $p+\mathrm{Pb}$ and $p+p$ cross sections are correlated. The uncertainty bands in the nPDF calculations reflect the variation of the nPDF parameters added in quadrature. This variation is a reflection of the precision of the data used to extract the nPDFs and is the leading uncertainty above $p_{T}\gtrsim 3$ GeV. At lower $p_{T}$, the scale uncertainties become very large but are not well defined because the factorization scale becomes smaller than the nPDF parametrization scale.
In general, the CGC predictions for single-inclusive hadrons tend to have somewhat steeper slopes as a function of $p_{T}$ and $y$ compared to the nPDF predictions. The trends associated with charged hadron production in the CGC approach are in somewhat better agreement with the LHCb data [324]. However, including the baryon contribution to charged hadron production results in a poorer description below $p_{T}\sim 10$ GeV using standard independent fragmentation functions [318]. The discrepancy with the data is probably most notable for the $p/\pi$ ratio which exhibits a bump at $p_{T}\sim 3$ GeV [326] that cannot be described by modifying the fragmentation functions. Indeed, the pion $p_{\rm T}$ slope is flatter [325] and better in line with the nPDF calculations.
Because the CGC and nPDF calculations both depend on the fragmentation functions, forward pion measurements differential in $p_{T}$ and $y$ would be more beneficial than inclusive charged hadrons alone. Based on Fig. 10, the nPDF and CGC calculations could be distinguished by systematic studies of the $p_{T}$ and $y$ slopes of the suppression factor at $p_{T}\gtrsim 3$ GeV. In $D$ meson studies this $p_{T}$ restriction could be relaxed because the finite charm quark mass allows the perturbative calculation to be continued to $p_{T}=0$. Existing LHCb measurements in the forward region [124, 127] could be studied from this perspective.
@@
In single inclusive hadron production (SIDIS), a final-state hadron with $p_{T}$ and $z$ could also distinguish between saturation and shadowing effects.
@@
Future SIDIS measurements at the EIC are known [327] to significantly reduce the uncertainties in the PDFs in $e+p$ collisions, particularly for radiative production of light sea quarks. Analogous measurements in $e+A$ collisions should therefore provide valuable information about sea quark relative to gluon shadowing of the nPDFs.
@@
At small $x$, the dominant parton-level production channel for SIDIS is $\gamma^{*}g\to q\bar{q}$ where the kinematics of one of the final state quarks is integrated over. SIDIS has the advantage of probing smaller values of $x$ for $t$-channel gluons compared to dihadron production, discussed in the previous section, because the gluon $x$ is comparable to Bjorken $x$ for $p_{T}\ll Q$.
@@
In fact, when $p_{T}\ll Q$, SIDIS also exhibits TMD factorization [328, 329, 294, 330] for the sea quarks: the CGC cross section can be expressed [331] as product of a hard factor for the $\gamma^{*}q\to q$ subprocess at scale $Q$ and a TMD sea quark distribution [332, 333, 334] resulting from $g\to q\overline{q}$ splitting in the nuclear target with the $q$ or $\overline{q}$ measured with final-state $p_{T}$. SIDIS can thus probe saturated sea quarks in the target wavefunction with $p_{T}\sim Q_{s}\ll Q$ [331]. An additional interesting regime of gluon saturation is the very forward limit [335], where $1-z\ll 1$, such that $(1-z)Q^{2}\sim Q_{s}^{2}$ with $Q\gg Q_{s}$. In this regime, the CGC predicts strong suppression in $R_{eA}(z)$ as $z\to 1$ [335].
@@
Due to the hierarchy of scales imposed in these situations, radiative corrections are dominated by Sudakov logarithms, similar to two-particle correlations, as previously discussed. The all-order exponentiation of these logarithms will also suppress the cross section, competing with the suppression caused by saturation. As for two-particle correlations, these corrections make the physical interpretation of the suppression more difficult. Theoretical frameworks including both suppression mechanisms, either within the TMD factorization formalism [297] or by coupling the CGC to a final-state parton shower [336, 303], are the only way to conclusively determine the role of nonlinear effects in future experimental data.
@@
In the absence of a specific hierarchy of scales imposed on the final state, the Sudakov effects disappear. However, the interesting regime for saturation is typically semi-hard scales where $p_{T}\sim Q\sim Q_{s}$ and $z\sim 0.5$. In this regime, no approximation is applicable and the all-twist CGC cross section must be employed. While LO estimates are simple to obtain because the SIDIS cross-section involves only dipoles, NLO predictions are challenging to compute numerically due to the complexity of the formulas [199, 195, 200]. A direct comparison to the leading-twist nPDF approach necessitates at least NLO precision. More developments are needed to evaluate CGC cross sections at full NLO.
@@
Single-inclusive jet production [337, 338, 339, 340], which has a reduced dependence on nonperturbative effects in the final state because no fragmentation functions are required, is related to SIDIS. At low $p_{T}$, $p_{T}\ll Q$, jets can probe the sea quark distribution, much like single inclusive hadron production [341]. Recently, jet distance measures have been developed to ensure TMD factorization beyond LO using jets [342, 341]. However, jet measurements at low $p_{T}$ remain challenging [343]. (At HERA, jets have been measured down to $4-5$ GeV in the Breit frame.) In addition, the phase space for low $p_{T}$ jets at high $Q^{2}$ and small $x$ shrinks drastically in the EIC kinematics. At moderate $x$, single-inclusive jet production at low $p_{T}$ is also sensitive to jet energy loss in CNM [344], see the discussion in the following section. Lastly, it is also possible to study saturation effects in jet SIDIS in the target fragmentation region of DIS in the Breit frame [345]. Furthermore, nucleon energy correlators have been proposed as a new way to study saturation dynamics in this kinematic regime [346].
@@ III.4.3 Diffractive vector meson production in $\gamma+A$
Figure 11: A compilation of exclusive $J/\psi$ production measurements in Pb+Pb UPCs at $\sqrt{s_{NN}}=5.02\,{\rm GeV}$ [135, 136, 347, 139, 348]. (Top) The curves represent NLO calculations with EPPS21 nPDFs using different factorization and renormalization scales $\mu$. The light blue band is the EPPS21 uncertainty at $\mu=2.39$ GeV. The NLO calculations are from Refs. [349, 350]. (Bottom) The calculations, taken from Ref. [351], represent different realizations of saturation-based calculations, depending on their description of the dipole scattering with the nucleus, are based on Refs. [352, 353, 354, 355, 356, 357].
@@
Diffractive vector meson production in $\gamma+A$ interactions serves as a complementary and particularly sensitive probe of gluon structure at small $x$ [358] relative to the observables discussed earlier. Unlike semi-inclusive observables, diffraction demands a color neutral exchange, which at LO is achieved by at least two-gluon exchange, making it more sensitive to saturation and shadowing effects. There are two kinds of reactions. In the first, coherent interactions, the nucleus remains intact and the average gluon distribution is probed. In the second, incoherent interactions are characterized by nuclear breakup and are sensitive to event-by-event fluctuations of the gluon densities [359, 360, 361, 358]. Diffractive $J/\psi$ production is particularly compelling because the charm mass provides a hard scale, allowing perturbative treatment, but the scale is also comparable to the saturation scale in $\gamma+A$ collisions.
@@
In the saturation formalism, this process is studied within the dipole picture, where, at LO, the photon splits into a $q\overline{q}$ pair [362] (and $q\overline{q}g$ at NLO [187]) which interacts with the nucleus through multiple coherent scatterings and then recombines into the vector meson. The saturation physics is encoded in the $q\overline{q}$ interaction with the nucleus, characterized by the dipole amplitude. The energy dependence of the dipole is governed by the BFKL equation in the linear regime and the BK/JIMWLK equation in the nonlinear regime.
@@
The cross section for diffractive vector meson production in $\gamma+A$ interactions has been studied as a function of the center of mass energy and rapidity of the produced vector meson in ultraperipheral collisions of heavy-ions at RHIC and the LHC. In the linear regime, governed by BFKL evolution, the cross section is predicted to grow as a power of the energy while saturation effects are expected to suppress this growth. The suppression is expected to be strong when the typical transverse size of the vector meson is larger than or comparable to the inverse of the saturation scale [363]. At the energies and rapidities reached in these colliders, the growth of the $J/\psi$ cross section in $\gamma+p$ collisions in UPCs is consistent with a power law because the typical transverse size of $J/\psi$ is smaller than the inverse of the saturation scale in the proton [364, 365]. On the other hand, strong nuclear suppression is seen in $J/\psi$ production in $\gamma+{\rm Pb}$ collisions at the LHC. The suppression is stronger than predicted by the current saturation-based approaches, see the lower panel of Fig. 11. These saturation-based predictions differ in their description of the dipole amplitude but are usually constrained by fits to $\gamma+p$ data. (The acronyms here refer to the models in the legend of the bottom panel of Fig. 11.) The CGC (MSS) calculation starts with initial conditions taken from an impact-parameter dependent MV model with the rapidity and energy evolution following the JIMWLK equation [352]. The b-BK (BCCM) model employs an impact-parameter dependent dipole amplitude that evolves according to impact-parameter dependent BK evolution [353]. The BGK-I (LM) [354], IPsat (LM) [355], GG-HS (CCK) [356], and IIM BG (GM) [357] calculations implement different phenomenological models of the impact parameter, rapidity and dipole size dependence of the dipole amplitude. For more details about the model implementation, see Ref. [351].
@@
Exclusive $J/\psi$ production in Pb+Pb collisions probes the nPDFs predominantly at $10^{-5}\lesssim x\lesssim 10^{-2}$ and scales similar to those in low $p_{T}$ open charm production [349, 350]. The coefficient functions are known up to NLO in strong coupling [366] when the $J/\psi$ is treated nonrelativistically. Extraction of the nPDFs, including testing dynamical modeling as in leading-twist shadowing [148, 367], based on UPC data is complicated by several factors. Convergence of the perturbative expansion has proven to be quite difficult due to coincidental cancellations between the LO and NLO gluon contributions. As a result, the calculated cross sections are rather sensitive to the factorization and renormalization scales, as shown in the upper panel of Fig. 11. Better stability seems to be achieved by including resummation of BFKL-type terms [368]. Other complications arise because generalized nuclear parton distributions enter the calculations. These GPDs can be approximated by nPDFs but the associated theoretical uncertainty is difficult to reliably assess and at least induces some degree of model dependence. Other theoretical uncertainties include the role of nonrelativistic corrections to the treatment of the $J/\psi$ wave function as well as theoretical modeling of the exclusivity condition. All in all, an accurate description of exclusive $J/\psi$ production in Pb+Pb UPCs is still a distant goal which hinders the determination of the dynamical effects of saturation. Currently there also appears to be tensions between the ALICE [135] and preliminary ATLAS [348] data, see the difference in the data at midrapidity in Fig. 11.
@@
Looking forward, studying the energy dependence in different colliding systems, such as $\gamma+{\rm O}$, as well as the production of lighter and heavier vector mesons such as $\phi$ and $\Upsilon$ could help to further discriminate among different approaches to the observed nuclear suppression and thus distinguish between saturation and nonperturbative shadowing. Moreover, the production ratios of excited vector meson states to the ground states, particularly $\psi$(2S) to $J/\psi$, offers additional discriminatory power. In Ref. [369] it was argued that the distinctive node in the $\psi$(2S) wavefunction enhances the sensitivity to nonlinear effects, providing a promising ameans to discriminate between linear shadowing and nonlinear saturation dynamics. Lastly, further measurements of incoherent reactions, where the nucleus breaks up, and which are sensitive to fluctuations, could also shed some light on small-$x$ dynamics [370].